AEY Inc is a company based in the U.S. assuming growth for them is set at 5%, the current risk free rate in the U.S. is 3% and market premium is 8%. AEY Inc. paid a dividend yesterday of $1.20 and has a Beta of 1.3, What is current price per share? If the Beta was 0.8 (all else remains constant) what would be new price per share?
Answer #1

r or required rate of return

=3%+1.3*8%
=13.40%

What is current price per share

=D1/(r-g)

=(1.20*(1+5%))/(13.40%-5%)
=15.00

new r or required rate of return

=3%+0.8*8%
=9.40%

What is newprice per share

=D1/(r-g)

=(1.20*(1+5%))/(9.40%-5%)
=28.64

the above is answer..